OpenStudy (anonymous):

There were 150 freshman and sophomores at the high school musical Reese and her friends attended. Freshman tickets were $2.00 each and sophomore tickets were $3.00 each. The receipts for freshman and sophomore tickets totaled $390. How many sophomores bought tickets?

OpenStudy (anonymous):

urgg why do you do this too me.. alright let me try

OpenStudy (anonymous):

Haha because you have to learn too :P

OpenStudy (anonymous):

true but its hard! do i plug in the numbers too 2F + 3S = 400?

OpenStudy (anonymous):

@DoomDude You want to help him figure the rest out ill brb!

OpenStudy (anonymous):

but isnt it supposed to equal 400?

OpenStudy (anonymous):

One sec i just confused myself, im going to write it in a different way!

OpenStudy (anonymous):

he easiest way to figure this out is to make two equations from the information given. First, we know the total number of freshman and sophomores that attended were 150. So: f + s = 150 Second, we know that the total profit made was $390 from $2 freshmen and $3 sophomore tickets. So: 2f + 3s = 390 So the two equations are: f + s = 150 2f + 3s = 390 Let's now isolate one variable in the first equation (I'll do it with f, since we want to find s; isolating s will give you an extra (though simple and easy) step to do): f = 150 - s Now plug this isolated variable into the second equation for "f": 2f + 3s = 390 2(150 - s) + 3s = 390 Now solve for s (which will give you the number of sophomores that bought tickets): 2(150 - s) + 3s = 390 (Simplified by distributing the 2) 300 - 2s + 3s = 390 (Simplify by adding -2s and 3s) 300 + s = 390 (Subtract 300 from both sides) s = 90 so it's 90.

OpenStudy (anonymous):

you sure? lol how do you know this so well?

OpenStudy (anonymous):

Well first off, ive learned this and secondly i have a geometry teacher online lol :) He's on my yahoo!

OpenStudy (anonymous):

lmao lucky! i got another question im about to put up.. i have online classes -.-

OpenStudy (anonymous):

Me too! Lol and okay